OpenStudy (anonymous):

Help Me<3 Part 1: Solve |2x+7| -3 > 10. You must show all work. Part 2: Use complete sentences to describe the graph in words.

OpenStudy (anonymous):

add 3 on both sides, try to get the absolute value by itself \[|2x+7|>13\]

OpenStudy (anonymous):

Now what 2 equations can you write from what @marird15 wrote. Think about what absolut value means

OpenStudy (anonymous):

2x+7>13 or -2x-7>-13? im not sure about the seond one

OpenStudy (anonymous):

you have it right except nothing happens to the 2x+7

OpenStudy (anonymous):

Do you see why?

OpenStudy (anonymous):

so only the 13 becomes a negative?

OpenStudy (anonymous):

ya because the expression is saying that (2x+7) = a positive or negative 13. because when the abs value is applied it becomes positive. If you apply the negative sign to the 2x+7 you are changing the equation

OpenStudy (anonymous):

I'd suggest working out the math for both just so you can see it

OpenStudy (anonymous):

if you do the math for your 2 equations you wrote above you'll see that they are the same equation. 2x+7=13 <-> -2x-7=-13 <-> -1(2x+7)=-1(13)

OpenStudy (anonymous):

Now you'll have to refresh me on something. I vaguely remember that you need to flip the < when you do what?